# Eigenvector decorrelation for random matrices

### v2

@@ 4.2. Zig step: Propagating bounds via the characteristic flow
For Hermitian $D_{j}\in\mathbf{C}^{N\times N}$ with $\langle D_{j}\rangle=0$, spectral parameters $z_{j}\in\mathbf{C}\setminus\mathbf{R}$, $j=1,2$, and fixed $T>0$ the characteristic flow is defined by the following ODEs (see also [32, (5.3)]):
@@ 4.2.1. Preliminaries on the characteristic flow and admissible control parameters
    [$\gamma$ is a lower bound on the stability operator] It holds that
@@
The proof of Proposition 4.5 and a sufficient condition for $D$ to satisfy (3.5) for $\mathbf{I}=\mathbf{R}$ are given in Appendix A.2.
As discussed around (4.1), during the proof of Theorem 3.2 we need to handle resolvent products of the form $GBGBG$. More precisely, let $D_{l}\in\mathbf{C}^{N\times N}$ be Hermitian deformations and $z_{l}\in\mathbf{C}\setminus\mathbf{R}$ for $l\in[3]$. Denote $G_{l}:=(W+D_{l}-z_{l})^{-1}$, $\nu_{l}:=(z_{l},D_{l})$ and $M_{l}:=M^{D_{l}}(z_{l})$. We define the deterministic approximation of $G_{1}B_{1}G_{2}B_{2}G_{3}$ by (see [27, Definition 4.1])
@@
$\ell(z_{1},z_{2}):=\eta_{1}\rho_{1}(z_{1})\wedge\eta_{1}\rho_{2}(z_{2}),\quad\text{where}\quad\eta_{j}=|\Im z_{j}|,\,\,\,\rho_{j}(z_{j})=\pi^{-1}|\langle\Im M^{D_{j}}(z_{j})\rangle|,\,\,\,j=1,2.\normalcolor$
@@
The proof of Proposition 4.6 is given in Appendix A.3.
@@ 4.2.2. Propagating local law bounds
Here, $B_{t-s}$ is a real symmetric ($\beta=1$) or complex Hermitian ($\beta=2$) matrix-valued Brownian motion with entries having variance equal to $(t-s)$ times those of a GOE/GUE matrix. Finally, we denote the resolvent of $W_{t}+D_{j,t}$ at $z_{j,t}$ by
@@ 4.3. Zag step: Removing Gaussian components via a GFT
(4.30) $\mathrm{d}W_{r}=-\frac{1}{2}W_{r}\mathrm{d}r+\frac{\mathrm{d}B_{r-s}}{\sqrt{N}}\quad\text{with initial condition}\quad W_{s}=W$
@@ 4.3.2. Conditional Gronwall estimates: general case
(4.41) $\left|\big{(}G_{1}B_{1}G_{2}B_{2}G_{1}\big{)}_{\bm{x}\bm{y}}\right|\leq\frac{\|B_{2}\|}{\eta}\sqrt{\big{(}G_{1}B_{2}\Im G_{2}B_{1}^{*}G_{1}^{*}\big{)}_{\bm{x}\bm{x}}\big{(}\Im G_{1}\big{)}_{\bm{y}\bm{y}}}\prec\frac{1}{\eta^{2}}$
@@ 4.3.3. Conditional Gronwall estimates: regular case
Fix $10\leq T<$ 1, and pick a spectral parameter $|z|\lesssim 1$ and a matrix $\lVert D\rVert\lesssim 1$. Then there exist initial conditions $z_{0},D_{0}$ such that the solutions $z_{t},D_{t}$ of (4.8), with initial conditions $z_{0},D_{0}$, satisfies $z_{T}=z$ and $D_{T}=D$. Additionally, we have $\mathrm{dist}(z_{0},\mathrm{supp}(\rho_{D_{0}}))\geq cT$, for some universal constant $c>0$.
@@
By Itô calculus, for any deterministic observables $R_{1},R_{2}\in\mathbf{C}^{N\times N}$, recalling (4.24), (4.8) and (4.23), we have take the real case. Probably the Setting(4.7) should also include the real case and introduce $\beta=1,2$ and then very soon in this proof we focus only on the complex case.]
(5.2) $\begin{split}\mathrm{d}\langle G_{1,t}R_{1}G_{2,t}R_{2}\rangle&=\mathrm{d}\mathcal{E}_{t}+\langle G_{1,t}R_{1}G_{2,t}R_{2}\rangle\mathrm{d}t+\langle G_{1,t}R_{1}G_{2,t}\rangle\langle G_{2,t}R_{2}G_{1,t}\rangle\mathrm{d}t\\ &\quad+\langle G_{1,t}-M_{1,t}\rangle\langle G_{1,t}^{2}R_{1}G_{2,t}R_{2}\rangle\mathrm{d}t+\langle G_{2,t}-M_{2,t}\rangle\langle G_{1,t}R_{1}G_{2,t}^{2}R_{2}\rangle\mathrm{d}t,\\ &\quad+\frac{\bm{1}(\beta=1)}{N}\bigg{[}\langle G_{1,t}^{\mathfrak{t}}G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}\rangle\mathrm{d}t+\langle G_{2,t}^{\mathfrak{t}}G_{2,t}R_{2}G_{1,t}R_{1}G_{2,t}\rangle\mathrm{d}t\\ &\qquad\qquad\qquad\quad+\langle(G_{1,t}R_{1}G_{2,t})^{\mathfrak{t}}G_{2,t}R_{2}G_{1,t}\rangle\mathrm{d}t\bigg{]}\,,\end{split}$
@@
The differential in (5.2) is complemented by the time derivative of the corresponding deterministic approximation (recall the shorthand notation $M_{12,t}^{R}$ from (3.9)) given in the next lemma. Its proof is completely analogous to [32, Lemma 5.5] and hence omitted.
@@
(5.7) $\begin{split}&\mathrm{Lin}_{t}=k(R_{1})\langle M_{12,t}^{R_{1}}\rangle g_{t}^{I,R_{2}}+k(R_{2})\langle M_{21,t}^{R_{2}}\rangle g_{t}^{R_{1},I},\\ &\mathrm{Err}_{t}=g_{t}^{I,R_{2}}g_{t}^{R_{1},I}+\langle G_{1,t}-M_{1,t}\rangle\langle G_{1,t}^{2}R_{1}G_{2,t}R_{2}\rangle+\langle G_{2,t}-M_{2,t}\rangle\langle G_{1,t}R_{1}G_{2,t}^{2}R_{2}\rangle\,,\end{split}$
@@
(5.9) $\begin{split}\tau^{R_{1},R_{2}}&:=\sup\{t\in[0,T]:\,\max_{s\in[0,t]}\max_{z_{j,0}\in\Omega^{j}_{0}}\alpha_{s}^{-1}\left|g_{s}^{R_{1},R_{2}}\right|\leq N^{2\xi_{k(R_{1},R_{2})}}\},\quad\normalcolor\\ \tau&:=\min\{\tau^{R_{1},R_{2}}:\,R_{1},R_{2}\in\mathfrak{S}\}\,,\quad\text{with}\quad\mathfrak{S}:=\{I,B_{1},B_{1}^{*},B_{2},B_{2}^{*}\}\,,\end{split}$
@@
We now state two important technical lemmas whose proofs are postponed to Section A.5 and after concluding the proof of (4.27a), respectively. Lemma 5.2 controls the propagator of (5.6).
@@
    For any $s,t\in[0,T]$, $s\leq t$, it holds $\beta_{s}\sim\beta_{t}+(t-s)$.
@@
In the following, we will consider (5.6) as a system of equations for $g^{R_{1},R_{2}}_{t}$, $R_{1},R_{2}\in\mathfrak{S}$. For each choice of $R_{1},R_{2}\in\mathfrak{S}$, we use the stochastic Gronwall argument from [38, Lemma 5.6] with (5.15) as an input to show that $\tau^{R_{1},R_{2}}>\tau$ unless $\tau^{R_{1},R_{2}}=T$. This would readily imply that $\tau=T$ with very high probability, i.e. (4.27a) holds.
Take any $R_{1},R_{2}\in\mathfrak{S}$ and denote $g_{s}:=g_{s}^{R_{1},R_{2}}$, $\xi:=\xi_{k(R_{1},R_{2})}$. Consider (5.15) for some $\zeta<\xi$. Due to the choice of $\xi_{j}$, $j\in[0,2]$ the rhs. of (5.15) is upper bounded by $\alpha_{t\wedge\tau}N^{\xi}$, where we ignored the irrelevant $\log N$ factor. Then [38, Lemma 5.6] with $d=1$ applied for the scalar equation (5.6) asserts that for any arbitrary small $\zeta>0$ and for any $t\geq 0$ we have
(5.16) $\sup_{0\leq s\leq t\wedge\tau}\left|g_{s}\right|^{2}\lesssim\left|g_{0}\right|^{2}+N^{2\xi+3\zeta}\alpha_{t\wedge\tau}^{2}+\int_{0}^{t\wedge\tau}\left(|g_{0}|^{2}+N^{2\xi+3\zeta}\alpha_{s}^{2}\right)f_{s}\exp\left(2\left(1+N^{-\zeta}\right)\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\mathrm{d}s.$
@@
$\exp\left(2N^{-\zeta}\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\leq\exp\left(CN^{-\zeta}\log N\right)\lesssim 1.$
@@
(5.17) $\sup_{0\leq s\leq t\wedge\tau}\left|g_{s}\right|^{2}\lesssim N^{2\xi+3\zeta}\alpha_{t\wedge\tau}^{2}+N^{2\xi+3\zeta}\int_{0}^{t\wedge\tau}\alpha_{s}^{2}f_{s}\exp\left(2\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\mathrm{d}s.$
@@
(5.19) $\int_{0}^{t\wedge\tau}\alpha_{s}^{2}f_{s}\exp\left(2\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\mathrm{d}s\leq\left(\frac{1}{N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau}}\right)^{2}\int_{0}^{t\wedge\tau}\left(\frac{\rho_{1,s}}{\eta_{1,s}}+\frac{\rho_{2,s}}{\eta_{2,s}}\right)\mathrm{d}s\lesssim\frac{\log N}{(N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau})^{2}}.$
To establish the upper bound of order $\log N/(N\ell_{t\wedge\tau}\gamma^{2}_{t\wedge\tau})$ for the lhs. of (5.18) we split the region of integration into two parts $[0,s_{*}]$ and $[s_{*},t\wedge\tau]$, where $s_{*}$ is defined as
@@
To prepare for the proof of Lemma 5.3 in the next proposition we show that the spectral domains $\Omega_{t}^{j}$ and $\Omega_{\kappa,t}^{j}$ (see (4.9) and (4.10)) for $t\in[0,T]$ and $j\in[2]$ satisfy the ray property. Informally this means that for every $z$ in these domains with $\Im z>0$ (resp. $\Im z<0$) the vertical ray going off toward $\Re z+\mathrm{i}\infty$ (resp. $\Re z-\mathrm{i}\infty$) is essentially contained in the domain. Since the result holds both for $j=1,2$, we will neglect $j$ in notations. The proof of Lemma 5.4 is given in Appendix A.5.
@@
    [Unrestricted spectral domains] For any $t\in[0,T]$, $z\in\Omega_{t}$ and $x\geq 0$ such that $|\Im z|+x\leq N^{100}$ it holds that $z+{\rm sgn}(\Im z)\mathrm{i}x\in\Omega_{t}$. That is, for $\Im z>0$ ($\Im z<0$) the vertical ray which starts at $z$, goes up (down) and leaves $\Omega_{t}$ only after reaching points with imaginary part larger than $N^{100}$ (smaller than $-N^{100}$).
@@
    [Bulk-restricted spectral domains] Fix a (small) $\kappa>0$. Then there exists $t_{*}\in[0,T]$ such that the previous part of the statement holds for $\Omega_{\kappa,t}$ for any $t\in[t_{*},T]$. Namely, For any $t\in[t_{*},T]$, $z\in\Omega_{\kappa,t}$ and $x\geq 0$ such that $|\Im z|+x\leq N^{100}$ it holds that $z+{\rm sgn}(\Im z)\mathrm{i}x\in\Omega_{\kappa,t}$. Moreover, $T-t_{*}\sim 1$ with implicit constants which depend only on $\kappa$ and $L$.
@@
(5.26) $\mathrm{QV}\big{[}g_{s}^{R_{1},R_{2}}\big{]}\lesssim\frac{1}{N^{2}}\left(\frac{1}{\eta_{1,s}^{2}}\langle\Im G_{1}R_{1}G_{2}R_{2}\Im G_{1}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle+\frac{1}{\eta_{2,s}^{2}}\langle\Im G_{2}R_{2}G_{1}R_{1}\Im G_{2}R_{1}^{*}G_{1}^{*}R_{2}^{*}\rangle\right).$
@@
$\frac{1}{N^{2}}\int_{0}^{t\wedge\tau}\frac{1}{\eta_{1,s}^{2}}\langle\Im G_{1}R_{1}G_{2}R_{2}\Im G_{1}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle\mathrm{d}s\lesssim\frac{1}{N^{2}}\int_{0}^{t\wedge\tau}\frac{\rho_{1,s}}{\eta_{1,s}^{3}\eta_{2,s}^{2}}\mathrm{d}s\lesssim\left(\frac{1}{N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau}}\right)^{2}.$
@@
In order to deal with absolute values of resolvents we employ the integral representation [34, Eq. (5.4)]:
@@
where we abbreviated $\zeta_{2,s,x}:=(\eta_{2,s}^{2}+x^{2})^{1/2}$. We now split the region of integration $[0,\infty)$ into two parts: $S_{1}$ corresponds to the regime $\zeta_{2,s,x}\leq N^{100}$ and $S_{2}$ is the complementary regime, i.e. $S_{2}:=[0,\infty)\setminus S_{1}$. Now, for any $x\in S_{1}$, by Lemma 5.4 it holds that $E_{2,s}+\mathrm{i}\zeta_{2,s,x}\in\Omega_{s}^{2}$. Thus we conclude
@@
it is sufficient to notice that $\alpha_{s}$ decreases along the flow and that by (3.6) and (4.12) we have
@@
(5.32) $|\langle G_{1}^{2}R_{1}G_{2}R_{2}\rangle|\leq\frac{\langle\Im G_{1}\rangle^{1/2}\langle\Im G_{1}R_{1}G_{2}R_{2}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle^{1/2}}{\eta_{1,s}}\leq\frac{\langle\Im G_{1}\rangle\lVert R_{1}G_{2}R_{2}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rVert^{1/2}}{\eta_{1,s}}\leq\frac{\rho_{1,s}}{\eta_{1,s}\eta_{2,s}}.$
@@
(5.37) $\mathrm{d}\big{(}G_{1,t}R_{1}G_{2,t}-M_{12,t}^{R_{1}}\big{)}_{\bm{vw}}=\left(1+(1-k(R_{1}))\langle M_{12,t}^{I}\rangle\right)\big{(}G_{1,t}R_{1}G_{2,t}-M_{12,t}^{R_{1}}\big{)}_{\bm{vw}}\mathrm{d}t+\mathrm{d}\mathcal{E}_{t}^{(2)}+\mathcal{F}_{t}^{(2)}\mathrm{d}t,\$
@@
(5.38) $\begin{split}&\mathrm{Lin}_{t}^{(2)}:=k(R_{1})\langle M_{12,t}^{R_{1}}\rangle\big{(}G_{1,t}G_{2,t}-M_{12,t}^{I}\big{)}_{\bm{vw}},\\ &\mathrm{Err}_{t}^{(2)}:=\big{\langle}G_{1,t}R_{1}G_{2,t}-M_{12,t}^{R_{1}}\big{\rangle}\left(G_{1,t}G_{2,t}\right)_{\bm{vw}}+\langle G_{1,t}-M_{1,t}\rangle\left(G_{1,t}^{2}R_{1}G_{2,t}\right)_{\bm{vw}}\\ &\qquad\qquad+\langle G_{2,t}-M_{2,t}\rangle\left(G_{1,t}R_{1}G_{2,t}^{2}\right)_{\bm{vw}}\,,\end{split}$
@@
(5.39) $\begin{split}&\mathrm{d}\big{(}G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{R_{1},R_{2}}\big{)}_{\bm{vw}}\\ &=\big{(}\tfrac{3}{2}+(2-k(R_{1},R_{2}))\langle M_{12,t}^{I}\rangle\big{)}\big{(}G_{1,t}R_{1}G_{2,t}R_{2}G_{1,t}-M_{121,t}^{R_{1},R_{2}}\big{)}_{\bm{vw}}\mathrm{d}t+\mathrm{d}\mathcal{E}_{t}^{(3)}+\mathcal{F}^{(3)}_{t}\mathrm{d}t\,,\end{split}$

### v3

@@ 4.2. Zig step: Propagating bounds via the characteristic flow
For Hermitian $D_{j}\in\mathbf{C}^{N\times N}$ with $\langle D_{j}\rangle=0$, spectral parameters $z_{j}\in\mathbf{C}\setminus\mathbf{R}$, $j=1,2$, and fixed $T>0$ the characteristic flow is defined by the following ODEs (see also [33, (5.3)]):
@@ 4.2.1. Preliminaries on the characteristic flow and admissible control parameters
    $\gamma$ is a lower bound on the stability operator] It holds that
@@
The proof of Proposition 4.5 and a sufficient condition for $D$ to satisfy (3.5) for $\mathbf{I}=\mathbf{R}$ are given in Section A.2 of the Supplementary Material [27].
As discussed around (4.1), during the proof of Theorem 3.2 we need to handle resolvent products of the form $GBGBG$. More precisely, let $D_{l}\in\mathbf{C}^{N\times N}$ be Hermitian deformations and $z_{l}\in\mathbf{C}\setminus\mathbf{R}$ for $l\in[3]$. Denote $G_{l}:=(W+D_{l}-z_{l})^{-1}$, $\nu_{l}:=(z_{l},D_{l})$ and $M_{l}:=M^{D_{l}}(z_{l})$. We define the deterministic approximation of $G_{1}B_{1}G_{2}B_{2}G_{3}$ by (see [28, Definition 4.1])
@@
$\ell(z_{1},z_{2}):=\eta_{1}\rho_{1}(z_{1})\wedge\eta_{2}\rho_{2}(z_{2}),\,\text{where}\,\,\eta_{j}=|\Im z_{j}|,\,\,\,\rho_{j}(z_{j})=\pi^{-1}|\langle\Im M^{D_{j}}(z_{j})\rangle|,\,\,j=1,2.\normalcolor$
@@
The proof of Proposition 4.6 is given in Section A.3 of the Supplementary Material [27].
@@ 4.2.2. Propagating local law bounds
Here, $B_{t-s}$ is a real symmetric ($\beta=1$) or complex Hermitian ($\beta=2$) matrix-valued Brownian motion with entries having variance equal to (ts)(t-s) times those of a GOE/GUE matrix. Finally, we denote the resolvent of $W_{t}+D_{j,t}$ at $z_{j,t}$ by
@@ 4.3. Zag step: Removing Gaussian components via a GFT
$\mathrm{d}W_{r}=-\frac{1}{2}W_{r}\mathrm{d}r+\frac{\mathrm{d}B_{r-s}}{\sqrt{N}}\quad\text{with initial condition}\quad W_{s}=W$
@@ 4.3.2. Conditional Gronwall estimates: general case
(4.40) $\left|\big(G_{1}B_{1}G_{2}B_{2}G_{1}\big)_{\bm{x}\bm{y}}\right|\leq\frac{\|B_{2}\|}{\eta}\sqrt{\big(G_{1}B_{1}\Im G_{2}B_{1}^{*}G_{1}^{*}\big)_{\bm{x}\bm{x}}\big(\Im G_{1}\big)_{\bm{y}\bm{y}}}\prec\frac{1}{\eta^{2}}$
@@ 4.3.3. Conditional Gronwall estimates: regular case
Fix $0\leq T<1$, and pick a spectral parameter $|z|\lesssim 1$ and a matrix $\lVert D\rVert\lesssim 1$. Then there exist initial conditions $z_{0},D_{0}$ such that the solutions $z_{t},D_{t}$ of (4.8), with initial conditions $z_{0},D_{0}$, satisfies $z_{T}=z$ and $D_{T}=D$. Additionally, we have $\mathrm{dist}(z_{0},\mathrm{supp}(\rho_{D_{0}}))\geq cT$, for some universal constant $c>0$.
@@
By Itô calculus, for any deterministic observables $R_{1},R_{2}\in\mathbf{C}^{N\times N}$, recalling (4.24), (4.8) and (4.23), we have
[…]
@@
The differential in (5.2) is complemented by the time derivative of the corresponding deterministic approximation (recall the shorthand notation $M_{12,t}^{R}$ from (3.9)) given in the next lemma. Its proof is completely analogous to [33, Lemma 5.5] and hence omitted.
@@
[…]
@@
[…]
@@
We now state two important technical lemmas whose proofs are postponed to Section A.5 of the Supplementary Material [27] and after concluding the proof of (4.27a), respectively. Lemma 5.2 controls the propagator of (5.6).
@@
    For any $\in[0,T], sts\leq t$, it holds $\beta_{s}\sim\beta_{t}+(t-s)$.
@@
In the following, we will consider (5.6) as a system of equations for $g^{R_{1},R_{2}}_{t}$, $R_{1},R_{2}\in\mathfrak{S}$. For each choice of $R_{1},R_{2}\in\mathfrak{S}$, we use the stochastic Gronwall argument from [39, Lemma 5.6] with (5.15) as an input to show that $\tau^{R_{1},R_{2}}>\tau$ unless $\tau^{R_{1},R_{2}}=T$. This would readily imply that $\tau=T$ with very high probability, i.e. (4.27a) holds.
Take any $R_{1},R_{2}\in\mathfrak{S}$ and denote $g_{s}:=g_{s}^{R_{1},R_{2}}$, $\xi:=\xi_{k(R_{1},R_{2})}$. Consider (5.15) for some $\zeta<\xi$. Due to the choice of $\xi_{j}$, $j\in[0,2]$ the rhs. of (5.15) is upper bounded by $\alpha_{t\wedge\tau}N^{\xi}$, where we ignored the irrelevant $\log N$ factor. Then [39, Lemma 5.6] with $d=1$ applied for the scalar equation (5.6) asserts that for any arbitrary small $\zeta>0$ and for any $t\geq 0$ we have
(5.16) $\begin{split}\sup_{0\leq s\leq t\wedge\tau}\left|g_{s}\right|^{2}\lesssim&\left|g_{0}\right|^{2}+N^{2\xi+3\zeta}\alpha_{t\wedge\tau}^{2}\\ &+\int_{0}^{t\wedge\tau}\left(|g_{0}|^{2}+N^{2\xi+3\zeta}\alpha_{s}^{2}\right)f_{s}\exp\left(2\left(1+N^{-\zeta}\right)\int_{s}^{t\wedge\tau}(1+f_{r})\mathrm{d}r\right)\mathrm{d}s.\end{split}$
@@
$\exp\left(2N^{-\zeta}\int_{s}^{t\wedge\tau}(1+f_{r})\mathrm{d}r\right)\leq\exp\left(CN^{-\zeta}\log N\right)\lesssim 1.$
@@
(5.17) $\sup_{0\leq s\leq t\wedge\tau}\left|g_{s}\right|^{2}\lesssim N^{2\xi+3\zeta}\alpha_{t\wedge\tau}^{2}+N^{2\xi+3\zeta}\int_{0}^{t\wedge\tau}\alpha_{s}^{2}f_{s}\exp\left(2\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\mathrm{d}s,$
where we incorporated the contribution from the integral of $1$ in the exponent into the implicit multiplicative constant.
@@
(5.19) $\begin{split}&\int_{0}^{t\wedge\tau}\alpha_{s}^{2}f_{s}\exp\left(2\int_{s}^{t\wedge\tau}f_{r}\mathrm{d}r\right)\mathrm{d}s\\ &\quad\leq\left(\frac{1}{N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau}}\right)^{2}\int_{0}^{t\wedge\tau}\left(\frac{\rho_{1,s}}{\eta_{1,s}}+\frac{\rho_{2,s}}{\eta_{2,s}}\right)\mathrm{d}s\lesssim\frac{\log N}{(N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau})^{2}}.\end{split}$
To establish the upper bound of order $\log N/(N\ell_{t\wedge\tau}\gamma^{2}_{t\wedge\tau})$ for the lhs. of (5.18) we split the region of integration into two parts $0,s_{*}]$ and $s_{*},t\wedge\tau]$, where $s_{*}$ is defined as
@@
To prepare for the proof of Lemma 5.3 in the next proposition we show that the spectral domains $\Omega_{t}^{j}$ and $\Omega_{\kappa,t}^{j}$ (see (4.9) and (4.10)) for $t\in[0,T]$ and $j\in[2]$ satisfy the ray property. Informally this means that for every $z$ in these domains with $\Im z>0$ (resp. $\Im z<0$) the vertical ray going off toward $\Re z+\mathrm{i}\infty$ (resp. $\Re z-\mathrm{i}\infty$) is essentially contained in the domain. Since the result holds both for $j=1,2$, we will neglect $j$ in notations. The proof of Lemma 5.4 is given in Section A.5 of the Supplementary Material [27].
@@
    [Unrestricted spectral domains] For any $t\in[0,T]$, $z\in\Omega_{t}$ and $x\geq 0$ such that $|\Im z|+x\leq N^{100}$ it holds that $z+{\mathrm{s}gn}(\Im z)\mathrm{i}x\in\Omega_{t}$. That is, for $\Im z>0$ ($\Im z<0$) the vertical ray which starts at $z$, goes up (down) and leaves $\Omega_{t}$ only after reaching points with imaginary part larger than $N^{100}$ (smaller than $N^{100}$).
@@
    [Bulk-restricted spectral domains] Fix a (small) $\kappa>0$. Then there exists $t_{*}\in[0,T]$ such that the previous part of the statement holds for $\Omega_{\kappa,t}$ for any $t\in[t_{*},T]$. Namely, For any $t\in[t_{*},T]$, $z\in\Omega_{\kappa,t}$ and $x\geq 0$ such that $|\Im z|+x\leq N^{100}$ it holds that $z+{\mathrm{s}gn}(\Im z)\mathrm{i}x\in\Omega_{\kappa,t}$. Moreover, $T-t_{*}\sim 1$ with implicit constants which depend only on $\kappa$ and $L$.
@@
(5.26) $\mathrm{QV}\big[g_{s}^{R_{1},R_{2}}\big]\lesssim\frac{1}{N^{2}}\left(\eta_{1,s}^{-2}\langle\Im G_{1}R_{1}G_{2}R_{2}\Im G_{1}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle+\eta_{2,s}^{-2}\langle\Im G_{2}R_{2}G_{1}R_{1}\Im G_{2}R_{1}^{*}G_{1}^{*}R_{2}^{*}\rangle\right).$
@@
$\frac{1}{N^{2}}\int_{0}^{t\wedge\tau}\!\!\eta_{1,s}^{-2}\langle\Im G_{1}R_{1}G_{2}R_{2}\Im G_{1}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle\mathrm{d}s\lesssim\frac{1}{N^{2}}\int_{0}^{t\wedge\tau}\frac{\rho_{1,s}}{\eta_{1,s}^{3}\eta_{2,s}^{2}}\mathrm{d}s\lesssim\left(\frac{1}{N\eta_{1,t\wedge\tau}\eta_{2,t\wedge\tau}}\right)^{2}.$
@@
In order to deal with absolute values of resolvents we employ the integral representation [35, Eq. (5.4)]:
@@
where we abbreviated $\zeta_{2,s,x}:=(\eta_{2,s}^{2}+x^{2})^{1/2}$. We now split the region of integration $0,\infty)$ into two parts: $S_{1}$ corresponds to the regime $\zeta_{2,s,x}\leq N^{100}$ and $S_{2}$ is the complementary regime, i.e. $S_{2}:=[0,\infty)\setminus S_{1}$. Now, for any $x\in S_{1}$, by Lemma 5.4 it holds that $E_{2,s}+\mathrm{i}\zeta_{2,s,x}\in\Omega_{s}^{2}$. Thus we conclude
@@
it is sufficient to notice that $\alpha_{s}$ increases along the flow and that by (3.6) and (4.12) we have
@@
(5.32) $\begin{split}|\langle G_{1}^{2}R_{1}G_{2}R_{2}\rangle|&\leq\frac{\langle\Im G_{1}\rangle^{1/2}\langle\Im G_{1}R_{1}G_{2}R_{2}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rangle^{1/2}}{\eta_{1,s}}\\ &\leq\frac{\langle\Im G_{1}\rangle\lVert R_{1}G_{2}R_{2}R_{2}^{*}G_{2}^{*}R_{1}^{*}\rVert^{1/2}}{\eta_{1,s}}\leq\frac{\rho_{1,s}}{\eta_{1,s}\eta_{2,s}}.\end{split}$
@@
[…]
@@
[…]
@@
[…]